Solve 3/5b +1 = 1/2b +0.4
Answer by stanbon(75887)

Solve (3/5)b +1 = (1/2)b + 4/10
-----------------------
Multiply thru by 10 to get:
6b + 10 = 5b + 4
b = -6
=============
Cheers,
Stan H.